If a man moves a small box that weighs 6N, 20 meters in 12 second, how much power
was used?
1
Expert's answer
2020-10-29T07:00:07-0400

If the force he applied was parallel to the horizontal ground, the work he did is


"W=Fd."

Now it is time to remember a popular joke: Can't find work? Multiply power by time!

So, the power is work over time:


"P=\frac{W}{t}=\frac{Fd}{t}=\frac{6\cdot20}{12}=10\text{ W}."